Sunland Company Ltd. publishes a monthly sports magazine, Fishing Preview. Subscriptions to the magazine cost $20 per year. During November 2017, Sunland sells 10,980 subscriptions for cash, beginning with the December issue. Sunland prepares financial statements quarterly and recognizes subscription revenue at the end of the quarter. The company uses the accounts Unearned Subscription Revenue and Subscription Revenue. The company has a December 31 year-end.

Prepare the entry in November for the receipt of the subscriptions. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

Nov.

enter an account title to record the receipt of the subscriptions in November

enter a debit amount

enter a credit amount

enter an account title to record the receipt of the subscriptions in November

enter a debit amount

enter a credit amount

eTextbook and Media

List of Accounts

Prepare the adjusting entry at December 31, 2017, to record subscription revenue in December 2017. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

Dec. 31

enter an account title for the adjusting journal entry on December 31, 2017

enter a debit amount

enter a credit amount

enter an account title for the adjusting journal entry on December 31, 2017

enter a debit amount

enter a credit amount

eTextbook and Media

List of Accounts

Prepare the adjusting entry at March 31, 2018, to record subscription revenue in the first quarter of 2018. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

Mar. 31

enter an account title for the adjusting journal entry on March 31, 2018

enter a debit amount

enter a credit amount

enter an account title for the adjusting journal entry on March 31, 2018
Answer #1
Date Account Titles and Explanation Debit Credit
Nov. Cash 219600 =10980*20
     Unearned Subscription Revenue 219600
Date Account Titles and Explanation Debit Credit
Dec. 31 Unearned Subscription Revenue 18300 =219600*1/12
       Subscription Revenue 18300
Date Account Titles and Explanation Debit Credit
Nov. Unearned Subscription Revenue 54900 =219600*3/12
       Subscription Revenue 54900